### v2

# Evaluating AI-Generated Essays with GRE Analytical Writing Assessment

Abstract
The recent revolutionary advance in generative AI enables the generation of realistic and coherent texts by large language models (LLMs). Despite many existing evaluation metrics on the quality of the generated texts, there is still a lack of rigorous assessment of how well LLMs perform in complex and demanding writing assessments. This study examines essays generated by ten leading LLMs for the analytical writing assessment of the Graduate Record Exam (GRE). We assessed these essays using both human raters and the e-rater® automated scoring engine as used in the GRE scoring pipeline. Notably, the top-performing Gemini and GPT-4o received an average score of 4.78 and 4.67, respectively, falling between ”generally thoughtful, well-developed analysis of the issue and conveys meaning clearly” and ”presents a competent analysis of the issue and conveys meaning with acceptable clarity” according to the GRE scoring guideline. We also evaluated the detection accuracy of these essays, with detectors trained on essays generated by the same and different LLMs.
Keywords: Large Language Models, Writing Assessment, GRE
1 Introduction
Recent research (W.-L. Chiang et al., 2023; OpenAI et al., 2024; Touvron, Lavril, et al., 2023; Touvron, Martin, et al., 2023) has shown that extensive pretraining on diverse corpora of human-authored texts and subsequent refinement via human feedback (Ouyang et al., 2022) have enabled large language models (LLMs) to produce coherent and contextually relevant writing products. Despite these significant advancements, AI-generated texts are not necessarily perfect. Several ongoing community efforts, such as the General Language Understanding Evaluation (GLUE) (A. Wang, 2018), its enhanced version SuperGLUE (A. Wang et al., 2019), and the more recent Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021), are evaluating the quality of LLM-generated texts to benchmark the progress. However, most of the aforementioned benchmarks are designed in multi-choice Question Answering format, and LLMs optimized for benchmark performances are less robust against simple dataset perturbation Alzahrani et al. (2024). Meanwhile, there are also efforts to use LLMs to evaluate LLM-generated texts automatically. For example, Fu et al. (2023) proposed an LLM-based evaluation protocol that interprets the model-predicted probability of specific token(s) as a quality score, while C.-H. Chiang & Lee (2023) and Liu et al. (2023) prompted LLMs to produce a Likert-style score ranging from 1 to 5 to assess LLM-generated texts.
From an assessment perspective, most of these evaluation metrics are based on relatively simple tasks, such as question answering, summarization, next-word prediction, and so on. There are also findings that suggest the LLM-based automated metrics could be unfaithful (P. Wang et al., 2023; Shen et al., 2023). As such, there is still a lack of rigorous evaluation of AI-generated writing products based on challenging writing tasks with a well-established scoring procedure. This study addresses this gap by leveraging the Graduate Record Examination (GRE) analytical writing assessment to evaluate AI-generated essays. The GRE analytical writing is well known for assessing the test-taker’s ability to think critically and communicate complex ideas effectively. It employs well-trained human raters and automated scoring engines, e-rater® (Attali & Burstein, 2006), to appraise written responses based on a multifaceted scoring framework (Educational Testing Service, n.d.-a). The assessment aims to gauge a candidate’s prowess in articulating and substantiating complex thoughts, constructing and assessing arguments, and maintaining a coherent and focused discourse throughout the essay.
We instructed several state-of-the-art LLMs to produce essays in response to GRE writing prompts. The selection of LLMs for this study encompasses the latest in AI technology as of mid-2024, including GPT-4o, Gemini, and Llama3-8b, as well as prior models dated back to mid-2023, including GPT-4, GPT-3.5 turbo, Google’s Bard, Llama, Vicuna, and Koala. We included these earlier models to explore potential trends that might emerge over time. All the generated essays were then evaluated by human raters and ETS’s automated scoring engine, e-rater®. By comparing the scores and language feature metrics of the generated essays, we aim to gain insights into AI-generated essays through a rigorous scoring framework set by the GRE program.
On the other hand, as AI technology advances, it becomes increasingly challenging to distinguish between human-written and AI-generated content (Tulchinskii et al., 2023; Weber-Wulff et al., 2023; Wu et al., 2023). This presents significant concerns, particularly in academic integrity and the preservation of authentic authorship. The detection accuracy of AI-generated texts is influenced by both the context and the contrast samples used. For large-scale writing assessments, AI-generated essays have been shown to be detectable with high accuracy (Hao & Fauss, 2024; Hao et al., 2024; Yan et al., 2023; Jiang et al., 2024a, b). However, as the number of high-performing LLMs continues to grow, it becomes increasingly important to benchmark the detectability of AI-generated essays across different LLMs and to assess the performance of detectors trained on various datasets.
This study aims to answer the following three research questions.
  • RQ1. What is the score distribution when LLM-generated essays are evaluated using the GRE analytical writing scoring framework?
  • RQ2. How do essays generated by different LLMs compare in terms of adherence to prompt-specified length, language features, textual similarity, and perplexity?
  • RQ3. How effective are the detectors in identifying LLM-generated essays using language and perplexity features when trained on essays produced by the same or different LLMs as those in the evaluation dataset?
2 Methods
2.1 Selection of LLMs
This study aims to evaluate the quality of essays generated by a spectrum of advanced LLMs, encompassing proprietary and open-source technologies, by May 20, 2024. Within the proprietary sector, we chose the top ones, including GPT-4o from OpenAI and Gemini-1.5 from Google. These closed-source models represent the forefront of AI research in the industry, offering a glimpse into the state-of-the-art language generation capabilities. In addition to these new models, we also included some older ones, such as GPT-4 and GPT-3.5 turbo from OpenAI and Bard from Google, which appeared in the middle of 2023.
On the other hand, we also included some leading open-source LLMs in our evaluation. Specifically, we considered the recent Llama-3 8B model, which has been the subject of extensive research since its release (Dubey et al., 2024). In addition, we experimented with some older models from the middle of 2023, including the Llama-1 13B (Touvron, Lavril, et al., 2023), as well as its two finetuned variants: Vicuna (W.-L. Chiang et al., 2023) and Koala (Geng et al., 2023), both of which underwent specialized optimization processes to further refine their language generation mechanisms. The 13-billion parameter version was selected for each of these earlier models. In addition to the Llama family of models from Meta, we included another popular LLM, Mistral-7B-Instruct-v0.1, from Mistral AI, which is comparable in size to the Llama-3 model. We include model details in Appendix B.
2.2 Prompts and Essay Generation
The GRE analytical writing consists of a 30-minute writing task, ”analyze an issue”, which requires examinees to ponder a topic, evaluate its intricacies, and formulate a well-reasoned argument buttressed by examples and explanations. All writing prompts of the GRE writing assessment are publicly available (ETS, n.d.-b). To alleviate the risk of biases introduced by an individual writing prompt, we randomly selected two writing prompts for this study. To prompt LLMs for generating the essays, we sent the original GRE writing prompt plus an instruction on length: ”Please keep the response to about 500 words.”, where the 500 words is the typical length of GRE essays. The resulting prompts are listed below.
  • Prompt 1: Governments should not fund any scientific research whose consequences are unclear. Write a response in which you discuss the extent to which you agree or disagree with the recommendation and explain your reasoning for the position you take. In developing and supporting your position, describe specific circumstances in which adopting the recommendation would or would not be advantageous and explain how these examples shape your position. Please keep the response to about 500 words.
  • Prompt 2: To understand the most important characteristics of a society, one must study its major cities. Write a response in which you discuss the extent to which you agree or disagree with the statement and explain your reasoning for the position you take. In developing and supporting your position, you should consider ways in which the statement might or might not hold true and explain how these considerations shape your position. Please keep the response to about 500 words
We generated 100 essays for each prompt from each LLM, leading to a total of 2,000 essays from the ten models. Each LLM comes with a set of hyperparameters that can be adjusted to control the text generation. Among these hyperparameters, temperature directly controls the variability of the generated texts. For comparison purposes, we set the temperature to 0.4 (This was used as the default value in earlier versions of the web-based ChatGPT to strike a balance between diversity and quality) for each model except Bard, where this parameter configuration was not accessible. Here, we caution that different LLMs may use varying scaling for their temperature, meaning that the same numerical value may affect the randomness and diversity of responses differently across models.
2.3 Evaluation and Detection
To address the first research question, we evaluated the essays based on the GRE scoring framework. The GRE program has a well-established scoring process involving well-trained human raters and an automated scoring engine, e-rater® (Educational Testing Service, n.d.-a). The human raters in this study were professional scoring experts in the GRE program, and they rated each essay with a score from 0 to 6 (see Appendix for more details). The e-rater® engine takes in an essay and outputs a final score as well as a set of language feature scores.
To address the second research question, we first checked how well the length of the generated essays adhered to our instructions in the prompt, as essay length is often positively correlated with essay scores in writing assessments. Then, we compared the pairwise similarity (semantic and verbatim) of generated essays, reflecting the diversity of the texts generated by the given LLMs. We then compared the e-rater® computed language features of these generated essays, such as grammar, usage, mechanics, styles, organization, development, and word complexity Attali & Burstein (2006); J. Chen et al. (2016). Table 1 below lists the definitions of a subset of the e-rater® language features we used in this study. Finally, we compared the perplexity of the generated essays, which characterizes how likely a sequence of words (essay) is generated from a given language model. Most LLMs used in this study do not allow the direct computation of perplexity. So, we chose GPT2 as a common language model to compute the perplexity of essays generated from all LLMs. As such, the perplexity here is a relative measure of how likely the essays have been generated by GPT2. Prior research has shown that perplexity based on a common language model is very informative for detecting AI-generated essays (e.g., Hao & Fauss (2024); Jiang et al. (2024a)).
To address the third research question, we randomly sampled 100 human-written essays corresponding to each of the above two prompts from the GRE writing test in 2021 when ChatGPT-like AIs were not available. These samples were mixed with essays generated by the LLMs discussed earlier. Based on the e-rater® engine-generated language features and perplexity-based features from these data, we developed machine learning classifiers to detect AI-generated essays and evaluate their performance.
Feature Description
Grammar A statistic based on the number of errors related to pronouns, run-ons, missing possessives, etc. The more negative the value, the more errors.
Mechanics A statistic based on the number of errors related to capitalization, punctuation, commas, hyphens, etc. The more negative the value, the more errors.
Style A statistic based on the number of errors related to repetition of words, inappropriate words, etc. The more negative the value, the more errors.
Usage A statistic based on the number of errors related to missing/wrong articles, nonstandard verbs, etc. The more negative the value, the more errors.
Organization The logarithm of the number of discourse units
Development The logarithm of the mean length of discourse units
Word Complexity The average length of words
Table 1: Language features from e-rater® engine output used in this study.
3 Results
3.1 Human Scores
We recruited expert human raters from the GRE program to score the 200 essays (100 from each of the two writing prompts) generated by each LLM. The raters followed the same scoring framework as in the GRE scoring operation and provided a score ranging from 0 to 6 for each essay (see Appendix for the definition of the scores). The source of the essay, i.e., written by humans or AI, has been deliberately hidden from the human raters, so they were unaware that some essays were generated by AI. The score distribution is show in Table 2. We observed that essays generated by the proprietary LLMs (e.g., GPT models and Gemini) generally received higher scores from human raters. Newer models in 2024 generally do better than the LLMs in 2023. The average scores are between 4 and 5, which correspond to ”presents a competent analysis of the issue and conveys meaning with acceptable clarity” and ”generally thoughtful, well-developed analysis of the issue and conveys meaning clearly” respectively according to the GRE scoring framework.
3.2 e-rater® Scores
In addition to human scores, we applied the e-rater® engine to score each AI-generated essay. The score distribution is listed in Table 2.
Type Scores GPT-4o Gemini Llama3-8b GPT-4 GPT-3.5 Bard Mistral Koala Vicuna Llama
(mid-2024) (mid-2023)
$faLock \faUnlock \faLock \faLock \faLock \faUnlock \faUnlock \faUnlock \faUnlock$
Human Rater 6 6.5% 6.5% 0.5% 2.5% 4.0% 0.5% 0.5% 0.5% 0.0% 0.0%
5 54.5% 64.5% 16.5% 48.0% 36.5% 12.5% 5.0% 2.5% 8.0% 0.0%
4 38.5% 29.0% 82.5% 49.5% 57.0% 75.5% 73.5% 72.5% 71.5% 1.5%
3 0.5% 0.0% 0.5% 0.0% 2.5% 11.5% 21.0% 24.0% 20.5% 20.5%
2 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.5% 0.0% 41.5%
1 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 32.0%
0 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 4.5%
Mean 4.67 4.78 4.17 4.53 4.42 4.02 3.85 3.79 3.88 1.82
e-rater® 6 31.5% 7.5% 1.5% 0.0% 3.5% 1.5% 1.0% 1.0% 0.5% 1.0%
5 68.5% 92.5% 93.0% 98.0% 93.5% 71.5% 52.5% 43.0% 54.0% 33.5%
4 0.0% 0.0% 5.5% 0.5% 3.0% 28.5% 44.5% 55.5% 45.5% 44.0%
3 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 2.0% 0.5% 0.0% 14.5%
2 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 5.0%
1 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.0%
0 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.0%
Mean 5.32 5.08 4.96 5.01 5.00 4.71 4.53 4.45 4.55 4.04
[…]
Compared with human scores, e-rater® tended to give higher scores to AI-generated essays, though the proprietary LLMs still outperform the open-source counterparts. This is unsurprising since e-rater® was developed and calibrated based on human-written essays and showed good agreement with human scores J. Chen et al. (2016). The current language features, such as grammar, usability, mechanics, and so on, are sufficient for scoring human-written essays but fall short of properly scoring AI-generated essays, which display almost impeccable language usage but with shallow meanings. In contrast, human raters adeptly recognize these shortcomings and assign more accurate scores to AI-generated essays.
3.3 Essay Length, Similarity, Language Features, and Perplexity
Figure 1: Distribution of the length of the generated essays. The red vertical dashed line indicates the mean essay length.
Essay Length Adherence
We introduced a length requirement, i.e., about 500 words, when prompting LLMs to generate essays. However, not all LLMs follow the instructions strictly. Figure 1 shows the distribution of the lengths of the generated essays in terms of the number of words. Overall, we observe that newer models exhibited smaller variances and better alignment with the 500-word instruction. For more recent models in 2024, GPT-4o and Gemini generated essays with lengths slightly greater than 500, while Llama3-8b generated an average of 460 words. For older models, close-sourced models performed better than open-sourced models (Mistral, Vicuna, Koala, and Llama) in following the length instruction. Essays generated by Llama displayed the largest length variability, indicating that the model struggled to consistently produce long and high-quality essays, which echoed the findings on essay quality.
Essay Similarity
The pairwise similarity of essays generated by the same LLM under the same prompt reflects consistency and variability under identical generation conditions (e.g., prompt, temperature). There are two commonly known types of similarities between texts: verbatim similarity, which describes how closely two texts match in exact wording or surface structure, and semantic similarity, which shows how similar two texts are in meaning, regardless of specific word choices. Comparing these two types of similarities for pairwise essays generated by LLMs under the same prompt could reveal how well the LLMs maintain meaning across different expressions of the same content.
For semantic similarity, we first convert each essay into a document vector using a popular neural embedding model jina-embedding-v3 model (Sturua et al., 2024). This embedding model was chosen based on its high performance on text similarity benchmarks of the MTEB English leaderboard [^1] https://huggingface.co/spaces/mteb/leaderboard and its superior performance on handling long documents. Once the essays are converted to vectors, we compute the cosine similarity between pairs of vectors corresponding to essays generated from the same LLM under the same prompt. Figure 2 shows the distribution of cosine similarity from two prompts. We observed that essays generated by the Llama model displayed the least similarity, comparable to human-written essays. Models such as GPT-4, Gemini, and GPT-4o tend to have distributions with higher mean cosine similarity, indicating that their generated essays are more semantically similar to each other, a sign of good alignment to the writing prompts. The distributions between the two prompts differ slightly, with some models demonstrating higher density and peaks for one prompt than another (i.e., GPT-4o has a higher density for prompt 1, while Llama3-8b obtains the highest mean similarity on prompt 2).
For verbatim similarity, we represent each essay as a trigram vector, a method proven effective in capturing textual similarities for test security applications (Choi et al., 2024). Based on the trigram vectors, we computed pairwise cosine similarities for each pair of essays from the same LLM and prompt. Figure 3 shows the distribution of the similarity. Overall, human essays displayed the lowest similarity, while different LLMs showed different levels of similar texts. Despite the essays from the same LLM and prompt showing a high level of semantic similarity, the verbatim similarity is much lower, though they are generally higher than that from human-written essays.
Figure 2: Semantic similarity between pairs of essays generated by the same LLM and prompt, measured using the cosine similarity of their embedding vectors.
Figure 3: Verbatim similarity between pairs of essays generated by the same LLM and prompt, measured using the cosine similarity of their trigram vector representation.
Language Features
The language features from e-rater® allow us to compare more detailed aspects of the generated essays across the LLMs. Table 3 shows the mean and its standard error of the language features. It’s important to note that grammar, mechanics, usage, and style features are negative by convention, as they represent errors—the closer the value is to 0, the fewer the errors. In contrast, organization, development, and word complexity features are positive, with higher values indicating better performance in these areas. To visualize the differences across LLM-generated and human-written essays, we apply MinMax scaling to bring all features into the same range (0 to 1) and show them in Figure 4. Based on the results, we have the following observations:
Humans are doing better in development but show weaker performance in grammar, mechanics, and usage. This suggests human writing is strong in structure and coherence but prone to technical errors compared to LLMs. The GPT family (particularly GPT-4o and GPT-4) shows strong performance in grammar, mechanics, and usage, with wide coverage on the radar plots, indicating an all-around high performance with fewer errors and more fluent text generation. The 2024 models generally perform better than the 2023 ones, suggesting steady progress in these LLMs.
Figure 4: Comparison of the language features (after MinMax scaling) of the essays from different LLMs and humans. The ones in blue colors are models in 2024 and the ones in green colors are those from 2023. The bigger the value, the better the essay quality in that dimension.
[…]
mid-2024
GPT-4o -0.075±0.001 -0.050±0.002 -0.035±0.002 -0.328±0.002 2.457±0.007 3.794±0.008 5.895±0.014
Gemini -0.080±0.001 -0.048±0.002 -0.059±0.002 -0.317±0.002 2.344±0.008 3.824±0.008 6.100±0.018
Llama3-8b -0.068±0.001 -0.014±0.002 -0.044±0.002 -0.392±0.002 2.233±0.008 3.836±0.007 5.547±0.015
mid-2023
GPT-4 -0.075±0.001 -0.022±0.002 -0.048±0.002 -0.356±0.002 2.359±0.007 3.762±0.007 5.610±0.018
GPT-3.5 -0.072±0.001 -0.013±0.001 -0.031±0.002 -0.378±0.002 2.343±0.011 3.865±0.010 5.339±0.015
Bard -0.082±0.001 -0.015±0.001 -0.038±0.002 -0.384±0.002 2.552±0.013 3.602±0.012 5.023±0.016
Mistral -0.082±0.001 -0.019±0.002 -0.037±0.002 -0.417±0.003 2.292±0.016 3.704±0.014 5.335±0.017
Koala -0.064±0.001 -0.010±0.001 -0.045±0.002 -0.420±0.002 2.208±0.011 3.773±0.010 5.147±0.017
Vicuna -0.067±0.001 -0.011±0.001 -0.034±0.002 -0.394±0.002 2.198±0.009 3.759±0.011 5.238±0.016
Llama -0.081±0.001 -0.036±0.004 -0.041±0.003 -0.344±0.008 1.830±0.046 4.428±0.052 4.648±0.037
Human -0.082±0.001 -0.103±0.002 -0.076±0.002 -0.333±0.002 2.114±0.012 4.290±0.013 5.111±0.015
Table 3: Mean and standard error of the mean corresponding to the key language features from different LLMs. $\uparrow$ means the negative values close to 0 are more favored.
Perplexity Distribution
Perplexity and perplexity-based features have been used to detect AI-generated essays. Specifically, it has been shown that essay perplexity alone can distinguish between ChatGPT-generated and human-written essays in high-stakes writing assessment (Hao & Fauss, 2024). In this study, we used GPT-2 as the underlying language model and adopted the implementation of perplexity by HuggingFace. [^2] https://huggingface.co/docs/transformers/perplexity The results are shown in Figure 5. The perplexity of human-written essays is much higher than that of AI-generated essays. Chronically, the newer models (GPT-4o, Gemini) showed slightly higher perplexity than older models, which suggests the essays generated by these models deviated further from the GPT-2 baseline model than the other LLMs. The perplexity of essays from Llama shows the largest variability, indicating that the essays may not be consistently generated.
Figure 5: Boxplot of the essay perplexity distribution for different models. The ones in blue colors are models in 2024 and the ones in green colors are those from 2023. The bigger the value, the better the essay quality in that dimension.
4 Detectability of AI-generated Essays
As the quality of essays generated by LLMs continues to improve, concerns are rising in education about students submitting AI-generated assignments, posing a significant challenge to academic integrity. The detection of AI-generated essays in large-scale assessments has been a widely discussed topic (Hao & Fauss, 2024; Jiang et al., 2024a; Yan et al., 2023), with most previous work focusing on detecting essays from a single LLM. However, with new and more powerful LLMs emerging every few months, it becomes increasingly difficult to identify which LLMs are being used. This highlights the importance of examining the detectability of AI-generated essays across multiple LLMs.
In particular, we consider two scenarios: within-model detectability and cross-model detectability. In the within-model scenario, we assess how well classifiers trained on essays generated by a specific LLM can detect essays produced by that same LLM. In the cross-model scenario, we examine how effectively a classifier trained on data from one LLM can detect essays generated by other LLMs. In both cases, the classifiers were developed using the seven language features in Table 3 and perplexity-based features (including the whole essay perplexity score, the mean, median, max, min, and 10, 20, 30, …, 90 percentiles of the sentence perplexity in each essay) respectively. We experimented with multiple popular classifiers, including XGBoost (T. Chen & Guestrin, 2016), Random Forest (Breiman, 2001), Gradient-Boosted Machine (Friedman, 2001), and Support Vector Machine (Cortes & Vapnik, 1995), and found that XGBoost consistently gave the best results. So, we only show the results of the XGBoost classifier.
For the within-model scenario, the results in Table 4 show that classifiers using perplexity features slightly outperformed those using the e-rater® language features, although both performed well overall. Notable exceptions include the GPT-4o and Gemini models, where language features are slightly more effective. This suggests that essays generated by these typical LLMs can be reliably detected in the context of GRE analytical writing as long as the LLM used is known. It’s important to note that these detection results apply to AI-generated essays without human revisions. If AI-generated text is revised by humans or mixed with human-written content, detection accuracy is expected to decrease Hao & Fauss (2024); Hao et al. (2024).
Feature Set LLM Accuracy Precision Recall F1 Score ROC AUC
Language GPT-4o 0.988 0.990 0.985 0.988 1.000
Perplexity 0.970 0.970 0.970 0.970 0.995
Language Gemini 0.995 0.995 0.995 0.995 0.997
Perplexity 0.910 0.939 0.880 0.908 0.960
Language Llama3-8b 0.972 0.970 0.975 0.972 0.998
Perplexity 0.997 0.995 1.000 0.998 0.997
Language GPT-4 0.975 0.976 0.975 0.975 0.997
Perplexity 0.997 0.995 1.000 0.998 1.000
Language GPT-3.5 0.957 0.956 0.960 0.957 0.995
Perplexity 0.997 1.000 0.995 0.997 0.997
Language Bard 0.960 0.958 0.965 0.960 0.997
Perplexity 0.997 1.000 0.995 0.997 0.997
Language Mistral 0.963 0.961 0.965 0.963 0.996
Perplexity 0.993 0.995 0.990 0.992 1.000
Language Koala 0.990 0.986 0.995 0.990 0.999
perplexity 0.997 0.995 1.000 0.998 0.997
Language Vicuna 0.985 0.976 0.995 0.985 1.000
Perplexity 0.997 0.995 1.000 0.998 0.997
Language Llama 0.948 0.955 0.940 0.947 0.988
Perplexity 0.997 1.000 0.995 0.997 0.997
Table 4: Performance of the XGBoost classifier with different feature sets: e-rater® language features, perplexity feature on the binary classification task. Results are based on the average of five-fold cross-validation.
The results for the cross-model case are shown in Figure 6. Overall, perplexity feature-based classifiers performed slightly better than language feature-based models when the training and testing data came from LLMs of the same year. They also excelled when trained on data from 2024 LLMs and tested on 2023 LLM-generated data. On the other hand, language feature-based classifiers did much better when they were trained on data from 2023 LLMs and tested on data from 2024 LLMs. Surprisingly, classifiers trained on GPT-4o data using perplexity-based features achieved very high accuracy in detecting essays from most LLMs, with the exception of essays generated by Gemini and the 2023 version of Llama. This indicates that perplexity features are generally effective across models but may struggle with certain newer or more distinct language patterns in specific LLMs.
(a) Language feature-based classifiers
(b) Perplexity feature-based classifiers
Figure 6: Detection accuracy when training the detector with one model and evaluating it on essays generated by different models. The left panel presents results using language-based features in the detector, while the right panel shows results using perplexity-based features. Results are based on the average of five-fold cross-validation.
5 Discussion
In this paper, we conducted a large-scale empirical evaluation of ten top LLMs by using GRE analytical writing assessment. Our goal is to establish benchmarks for the quality of essays generated by AI in cases that require high-level analytical thinking skills. Our results indicate that proprietary LLMs, such as GPTs and Gemini, consistently produced higher-scoring essays compared to open-source models, aligning with existing evaluations that show proprietary models outperforming their open-source counterparts. Additionally, we observed a notable improvement in essay scores from 2024 models compared to 2023 models, reflecting the continuous efforts within the AI community to create increasingly capable models for generating higher-quality text.
Further analysis of the language features computed by the e-rater® engine revealed that AI-generated essays better-controlled grammar, usage, and mechanics errors than human-written essay baseline. All AI-generated essays show a lower perplexity than the baseline of human-written essays. These findings provide information for improving future AI text generation, emphasizing crafting more meaningful texts alongside its already high-quality linguistic characteristics.
Additionally, we have devised several machine learning classifiers, using language features from e-rater® and perplexity features to distinguish between human-written and AI-generated essays. The performance of all these classifiers is promising, suggesting that AI-generated essays responding to GRE analytical writing prompts can be reliably identified. However, these results apply mainly when no further revisions have been made to the AI-generated essays. Nevertheless, these findings establish a benchmark for the detection accuracy of such essays within the context of GRE analytical writing, and the findings are consistent with findings from other writing assessments (Yan et al., 2023; Hao & Fauss, 2024).
To summarize, this study has evaluated the essays generated by LLMs using the GRE analytical writing framework and benchmarked various properties under controlled conditions. We hope the findings will help the assessment community understand the status of AI-generated essays and the potential implications of writing assessments. Meanwhile, we hope the findings can also help the LLM community think about ways to enhance the quality of model outputs through improved training and fine-tuning techniques.
Limitations
The current study has several limitations. Firstly, LLMs are constantly evolving, and it is challenging to enumerate all available advanced models. Whether and how the findings of this paper will change as LLMs change needs follow-up investigation, which we will report in future work.
Secondly, it is crucial to note that the quality of generated texts heavily hinges on the provided prompts. In our study, we used the original prompts from the analytical writing assessment, only adding a length requirement. However, we acknowledge that modifying prompts in several ways, such as instructing the model to write as if it were a six-year-old, can dramatically alter the quality of the generated essay. As there are infinite ways to tweak the prompts, our choice is motivated by the idea of modifying the prompts as little as possible to simplify and obtain baseline results for benchmarking purposes. In future work, we will explore the impact of prompt tweaking on the essay quality.
Lastly, the variability of the generated essay is controlled by hyper-parameters, such as the temperature of the LLMs. Different LLMs may have different baseline settings for these hyper-parameters, which leads to varying degrees of variability in generated essays even when using the same hyper-parameter values. Therefore, a systematic investigation of the quality of essays for different hyper-parameters is imperative, and we will report the findings in future work.
Acknowledgment
YZ thank the ETS Research Summer Internship Program to fund the project. We also thank ETS Research Institute for funding the project.
Appendix A GRE Scoring Framework
The scoring framework of GRE “analyze an issue” writing task (Retrieved from GRE General Test Analytical Writing Scoring (ets.org))
Score 6: Outstanding
A 6 response presents a cogent, well-articulated analysis of the issue and conveys meaning skillfully. A typical response in this category:
  • Articulates a clear and insightful position on the issue in accordance with the assigned task.
  • Develops the position fully with compelling reasons and/or persuasive examples.
  • Sustains a well-focused, well-organized analysis, connecting ideas logically.
  • Conveys ideas fluently and precisely, using effective vocabulary and sentence variety.
  • Demonstrates superior facility with the conventions of standard written English (i.e., grammar, usage, and mechanics), but may have minor errors
Score 5: Strong
A 5 response presents a generally thoughtful, well-developed analysis of the issue and conveys meaning clearly. A typical response in this category:
  • Presents a clear and well-considered position on the issue in accordance with the assigned task.
  • Develops the position with logically sound reasons and/or well-chosen examples.
  • Is focused and generally well organized, connecting ideas appropriately.
  • Conveys ideas clearly and well, using appropriate vocabulary and sentence variety.
  • Demonstrates facility with the conventions of standard written English but may have minor errors
Score 4: Adequate
A 4 response presents a competent analysis of the issue and conveys meaning with acceptable clarity. A typical response in this category:
  • Presents a clear position on the issue in accordance with the assigned task.
  • Develops the position with relevant reasons and/or examples.
  • Is adequately focused and organized.
  • Demonstrates sufficient control of language to express ideas with acceptable clarity.
  • Generally demonstrates control of the conventions of standard written English, but may have some errors
Score 3: Limited
A 3 response demonstrates some competence in addressing the specific task directions, in analyzing the issue and in conveying meaning, but is obviously flawed. A typical response in this category exhibits one or more of the following characteristics:
  • Is vague or limited in addressing the specific task directions and in presenting or developing a position on the issue or both.
  • Is weak in the use of relevant reasons or examples or relies largely on unsupported claims.
  • It is limited in focus and/or organization.
  • Has problems in language and sentence structure that result in a lack of clarity.
  • Contains occasional major errors or frequent minor errors in grammar, usage or mechanics that can interfere with meaning
Score 2: Seriously Flawed
A 2 response largely disregards the specific task directions and/or demonstrates serious weaknesses in analytical writing. A typical response in this category exhibits one or more of the following characteristics:
  • Is unclear or seriously limited in addressing the specific task directions and in presenting or developing a position on the issue or both.
  • Provides few, if any, relevant reasons or examples in support of its claims
  • Is poorly focused and/or poorly organized.
  • Has serious problems in language and sentence structure that frequently interfere with meaning.
  • Contains serious errors in grammar, usage or mechanics that frequently obscure meaning
Score 1: Fundamentally Deficient
A 1 response demonstrates fundamental deficiencies in analytical writing. A typical response in this category exhibits one or more of the following characteristics:
  • Provides little or no evidence of understanding the issue.
  • Provides little or no evidence of the ability to develop an organized response (e.g., is disorganized and/or extremely brief)
  • Has severe problems in language and sentence structure that persistently interfere with meaning.
  • Contains pervasive errors in grammar, usage or mechanics that result in incoherence
Score 0: Fundamentally Deficient
Off topic (i.e., provides no evidence of an attempt to address the assigned topic), is in a foreign language, merely copies the topic, consists of only keystroke characters or is illegible or nonverbal.
Appendix B LLM Model Details
Below, we listed the details of LLMs use used for essay generation.
LLM Version URL
GPT-4o 2024-05-13 https://openai.com/index/hello-gpt-4o/
Gemini (1.5 Pro) 2024-05-14 https://gemini.google.com
Llama3-8B-Instruct 2024-04-18 https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
GPT-4 2024-05-13 https://platform.openai.com/docs/models
GPT3.5-turbo 2023-06-13 https://platform.openai.com/docs/models
Bard 2023-06-07 (now upgraded to Gemini)
Mistral 2023-09-27 https://mistral.ai/news/announcing-mistral-7b/
Vicuna-13B 2023-03-30 https://lmsys.org/blog/2023-03-30-vicuna/
Koala-13B 2023-04-03 https://bair.berkeley.edu/blog/2023/04/03/koala/
Llama1-13B 2023-02-24 https://ai.meta.com/blog/large-language-model-llama-meta-ai/
Table 5: LLMs used for our experiments.